OpenStudy (anonymous):

MEDAL GIVAWAY! Andrew salary is 1200 a month plus 4% commission on all his sales. What must the amount of his sales be to earn at least 1600 each month?

OpenStudy (anonymous):

1600 - 1200 = 400 -- he needs to earn at least 400 more dollars. 4% of his sales = 400 .04x = 400 x = 400/.04 x = 10,000 check...1200 + (.04)(10,000) = 1600 1200 + 400 = 1600 1600 = 1600 (correct) He would have to sell at least $10,000 each month to earn $1600. Does that make sense to you ?

OpenStudy (anonymous):

Yes, thank you!